OpenStudy (anonymous):

What is: 2m-nx=x+4. and we are supposed to be solving for x.

OpenStudy (anonymous):

2m=x+4+nx 2m-4=x+nx 2m-4=(1+n) 2m-4/1+n x=2m-4/1+n

OpenStudy (anonymous):

Thank you

OpenStudy (triciaal):

do you understand the steps to be used?